# Smoothed Energy Guidance: Guiding Diffusion Models with Reduced Energy Curvature of Attention

### v1

@@ 3.2 Analysis of the energy landscape
From a probabilistic perspective, this process can be thought of as maximizing the likelihood of the attention weights in terms of the Boltzmann distribution conditioned on a given configuration, i.e., the feature map. Blurring the attention weights diminishes this likelihood as shown in Lemma 3.2, and also reduces the curvature of the distribution as shown in Theorem 3.1.
@@ 4 Discussion on related work
Tackling the caveats of CFG, unconditional approaches such as self-attention guidance (SAG) [17] and perturbed attention guidance (PAG) [1] have been proposed. SAG selectively blurs images with the mask obtained from the attention map and guides the generation process given the prediction. This indirect approach causes saturation and noisy images when given a large guidance scale, leading to the selection of a guidance scale less than or equal to $1$. PAG guides images using prediction with identity attention, where the attention map is an identity matrix. However, the reliance on heuristics to make perturbed predictions results in unintended side effects. As an example of the side effects of replacing the attention map with identity attention, PAG changes the visual structure and color distribution of an image, as evidenced in Figs. 5, 7, and 8.
@@ 5.3 Controlling image generation with the standard deviation
In Figs. 3, 4, 9, 10, and 13, we display sampling results conditioned on text, Canny, and depth map. Using text (Fig. 3), the vanilla SDXL without CFG is unable to generate high-quality images and produces noisy results. Canny and depth map conditioning on SDXL (Fig. 4, 9, and 10) is achieved through ControlNet [51], trained on such maps. The results show that SEG enhances the quality and fidelity of the generated images while preserving the textual and structural information provided by the conditioning inputs. Notably, as $\sigma$ increases, the generated images exhibit improved definition and quality without introducing significant artifacts or deviations from the original condition. The combination with higher CFG scales is shown in Figs. 1418.
@@ 5.4 Comparison with previous methods
Since the results are visually favorable when we use $\sigma=10$ and $\sigma\to\infty$, and they are the best in terms of CLIP score and FID, respectively, we adopt those configurations for comparison of unconditional guidance methods. The results are presented in Figs. 5, 7, 8, and Table 1. Notably, our method achieves better image quality in terms of FID, while remaining similar to the original output of vanilla SDXL as measured by LPIPS, implying a Pareto improvement.
@@ Appendix A Full proofs
Assuming $\xi(\tilde{\mathbf{a}})_{i}\xi(\tilde{\mathbf{a}})_{j}\approx 0$ and $\xi({\mathbf{a}})_{i}\xi({\mathbf{a}})_{j}\approx 0$ for all $i$ and $j$, which is a reasonable assumption when the number of token is large, the non-diagonal elements of the Hessians approximate to 00 and the diagonal elements dominate. Therefore, the determinants of the Hessian matrices are approximated as the product of the dominant terms:
@@ Appendix C Additional qualitative results
Figs. 7 and 8 provide a qualitative comparison of SEG against vanilla SDXL [35], Self-Attention Guidance (SAG) [17], and Perturbed Attention Guidance (PAG) [1]. These comparisons highlight the superior performance of SEG in terms of image quality, coherence, and adherence to the given prompts. SEG consistently produces sharper details, more realistic textures, and better overall composition compared to the other methods.
@@
Figs. 9 and 10 showcase the application of SEG in conjunction with ControlNet [51] for conditional image generation. These results illustrate how SEG can enhance the quality and coherence of generated images while maintaining fidelity to the provided control signals. The images demonstrate improved detail, texture, and overall visual appeal compared to standard ControlNet outputs without prompts.
@@
Figure 14: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a friendly robot helping an old lady cross the street". Without causing saturation or significant changes in the general structure, SEG improves the image quality.
Figure 15: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a skateboarding turtle zooming through a mini city made of Legos".
Figure 16: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a group of puppies playing soccer with a ball of yarn".
Figure 17: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a baby elephant learning to paint with its trunk in an art studio".
Figure 18: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a family of teddy bears having a barbecue in their backyard".

### v2

@@ 3.2 Analysis of the energy landscape
From a probabilistic perspective, the energy is associated with the likelihood of the attention weights in terms of the Boltzmann distribution conditioned on a given configuration, i.e., the feature map. Blurring the attention weights diminishes this likelihood as shown in Lemma 3.2, and also reduces the curvature of the distribution as shown in Theorem 3.1.
@@ 4 Discussion on related work
Tackling the caveats of CFG, unconditional approaches such as self-attention guidance (SAG) [17] and perturbed attention guidance (PAG) [1] have been proposed. SAG selectively blurs images with the mask obtained from the attention map and guides the generation process given the prediction. This indirect approach causes saturation and noisy images when given a large guidance scale, leading to the selection of a guidance scale less than or equal to $1$. PAG guides images using prediction with identity attention, where the attention map is an identity matrix. However, the reliance on heuristics to make perturbed predictions results in unintended side effects. As an example of the side effects of replacing the attention map with identity attention, PAG changes the visual structure and color distribution of an image, as evidenced in Figs. 5, 8, and 9.
@@ 5.3 Controlling image generation with the standard deviation
In Figs. 3, 4, 10, 11, and 14, we display sampling results conditioned on text, Canny, and depth map. Using text (Fig. 3), the vanilla SDXL without CFG is unable to generate high-quality images and produces noisy results. Canny and depth map conditioning on SDXL (Fig. 4, 10, and 11) is achieved through ControlNet [51], trained on such maps. The results show that SEG enhances the quality and fidelity of the generated images while preserving the textual and structural information provided by the conditioning inputs. Notably, as $\sigma$ increases, the generated images exhibit improved definition and quality without introducing significant artifacts or deviations from the original condition. The combination with higher CFG scales is shown in Figs. 1519.
@@ 5.4 Comparison with previous methods
Since the results are visually favorable when we use $\sigma=10$ and $\sigma\to\infty$, and they are the best in terms of CLIP score and FID, respectively, we adopt those configurations for comparison of unconditional guidance methods. The results are presented in Figs. 5, 8, 9, and Table 1. Notably, our method achieves better image quality in terms of FID, while remaining similar to the original output of vanilla SDXL as measured by LPIPS, implying a Pareto improvement.
@@ Conclusion
Acknowledgements
I would like to express my gratitude to Yong-Hyun Park, Junha Hyung, and Donghoon Ahn for their valuable feedback and insights. Their thoughtful comments and suggestions have been instrumental in improving this work.
@@ Appendix A Full proofs
Assuming $\xi(\tilde{\mathbf{a}})_{i}\xi(\tilde{\mathbf{a}})_{j}\approx 0$ and $\xi({\mathbf{a}})_{i}\xi({\mathbf{a}})_{j}\approx 0$ for all $i$ and $j$, which is a reasonable assumption when the number of token is large and the softmax values get small, the non-diagonal elements of the Hessians approximate to 00 and the diagonal elements dominate. Therefore, the determinants of the Hessian matrices are approximated as the product of the dominant terms:
@@ Appendix C Additional qualitative results
Figs. 8 and 9 provide a qualitative comparison of SEG against vanilla SDXL [35], Self-Attention Guidance (SAG) [17], and Perturbed Attention Guidance (PAG) [1]. These comparisons highlight the superior performance of SEG in terms of image quality, coherence, and adherence to the given prompts. SEG consistently produces sharper details, more realistic textures, and better overall composition compared to the other methods.
@@
Figs. 10 and 11 showcase the application of SEG in conjunction with ControlNet [51] for conditional image generation. These results illustrate how SEG can enhance the quality and coherence of generated images while maintaining fidelity to the provided control signals. The images demonstrate improved detail, texture, and overall visual appeal compared to standard ControlNet outputs without prompts.
@@
Ablation study
Fig. 20 displays a visual example of unconditional generation with controlled $\gamma_{\mathrm{seg}}$ and $\sigma$. Consistent with results in Sec. 5.5, controlling image quality with $\sigma$ has fewer side effects than controlling with $\gamma_{\mathrm{seg}}$.
Appendix D Pipeline figure
The overall pipeline and conceptual framework of SEG are presented in Fig. 7. Fig. 7 (a) and Fig. 7 (b) depict the original sampling process and the modified sampling process with smoothed energy, respectively. Fig. 7 (c) illustrates the the final prediction (the red arrow) with the guidance scale.
Figure 7: Pipeline of SEG. (a) Original sampling process, self-attention weights, and the corresponding energy landscape. (b) Our modified sampling process with blurred queries where $\sigma\in(0,\infty)$, inducing blurred attention weights and the corresponding smoothed energy landscape. (c) A conceptual figure of $\gamma_{\mathrm{seg}}$. Note that since the guidance linearly extrapolates predictions from (a) and (b), a high guidance scale causes samples to be out of the manifold.
@@ Appendix D Pipeline figure
Figure 15: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a friendly robot helping an old lady cross the street." Without causing saturation or significant changes in the general structure, SEG improves the image quality.
Figure 16: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a skateboarding turtle zooming through a mini city made of Legos."
Figure 17: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a group of puppies playing soccer with a ball of yarn."
Figure 18: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a family of teddy bears having a barbecue in their backyard."
Figure 19: Experiment on the combination of SEG and CFG. $\gamma_{\text{seg}}$ is fixed to $3.0$. The prompt is "a baby elephant learning to paint with its trunk in an art studio."
Figure 20: Unconditional generation result with controlled $\gamma_{\mathrm{seg}}$ and $\sigma$.